OpenStudy (anonymous):

simplify 7 sqrt of 28

OpenStudy (anonymous):

sqrt(28) = 2sqrt(7) therefore 7sqrt(28) = 7(2sqrt(7)) = 14sqrt(7)

OpenStudy (precal):

\[7\sqrt{28}\] \[7\sqrt{4*7}\] \[7*2\sqrt{7}\] \[14\sqrt{7}\]